OpenStudy (anonymous):

what is the name of this conic? : 4x^2 + 4y^2 + 16x - 24y + 8 = 0

myininaya (myininaya):

Circle

OpenStudy (anonymous):

thanks

myininaya (myininaya):

Since the coefficients of x^2 and y^2 are the same

OpenStudy (anonymous):

ok, and is this an ellipse: x^2 + 9y^2 - 14x + 36y + 49 = 0

OpenStudy (anonymous):

damn, myininaya you are GOOD at maths

myininaya (myininaya):

it is an ellipse since the coefficients are not the same but both positive

OpenStudy (anonymous):

ok thank you

myininaya (myininaya):

whatever andruis Apparently I'm not that good since I can't figure out your problem

myininaya (myininaya):

And there was no math involved here lol

OpenStudy (anonymous):

:D, still, you are like 100 times better than me

OpenStudy (anonymous):

how would you go about generating an equation for a parabola with diretrix x=6 and vertex (3,8) ??

myininaya (myininaya):

brb i have to pee but the equation of parabola is f(x)=a(x-h)^2+k where (h,k) here =(3,8). So now all we have to find is a

OpenStudy (anonymous):

whoa dude your aewsome

OpenStudy (anonymous):

awesome* lol

myininaya (myininaya):

honestly i never heard of a diretrix

myininaya (myininaya):

i have to look it up

OpenStudy (anonymous):

its "a fixed line" having to deal with a parabola, like the focus is a " fixed point"

myininaya (myininaya):

ok. The equation y^2=4px is a parabola with focus (p,0) and directrix x=-p so...

myininaya (myininaya):

omg wait

OpenStudy (anonymous):

ok

myininaya (myininaya):

i think im fixing to figure something out

OpenStudy (anonymous):

thanks ive been out of school for weeks cuz i broke my arm my collar bone and my back lol so i havnt learned any of this

myininaya (myininaya):

I think a is -1/24 since p=1/4a and p=-6 since the directrix is x=6

myininaya (myininaya):

so I think the parabola is f(x)=-1/24(x-3)^2+8

OpenStudy (anonymous):

thats great man i appreciate it

myininaya (myininaya):

let me rewrite this so there is no confusion f(x)=-1*(x-3)^2/24+8

myininaya (myininaya):

i'm not a man lol

OpenStudy (anonymous):

oops my bad

myininaya (myininaya):

I could be wrong though. you should check with someone else

OpenStudy (anonymous):

alright well thanks anyway